OpenStudy (anonymous):

Line L passes through points (-2,1), (-4,-5). Find the equations of line L in the form y=mx + c.

OpenStudy (jango_in_dtown):

y-y1=(y2-y1)/(x2-x1)(x-x1) put x1=-2 y1=1 x2=-4 and y2=-5 then you have y-1=2(x+2) this imply y=2x+5

OpenStudy (jango_in_dtown):

@kunal28k

OpenStudy (anonymous):

When I minused y2 and y1 I got -6 and therefore my equation read y=3x-2 ?

OpenStudy (anonymous):

@ybarrap help

OpenStudy (ybarrap):

You are looking for m and c in this equation: $$ y=mx+c $$ To get it, use two points. Using point (-2,1) we have that x=-2 and y=1: $$ 1=-2m+c $$ Using point (-4,-5) we have that x=-4 and y=5: $$ 5=-4m+c $$ So now we have two equations and two unknowns, m and c: $$ 1=-2m+c\\ 5=-4m+c $$ Solve this for m and c and substitute into $$ y=mx+c $$ And you are done. You will have an equation y in terms of x. Make sense?

OpenStudy (anonymous):

Great thank you very much for your help. This just made my life so much easier.